OpenStudy (petlovers123):

Which of the following demonstrates the Commutative Property of Multiplication? 2(6a − 3) = (6a − 3) ⋅ 2 2(6a − 3) = 12a − 6 12a − 6 = (6a − 3) ⋅ 2 (2 ⋅ 6a) − 3 = 2(6a − 3) Which pair of expressions is equivalent using the Associative Property of Multiplication? 6(4a ⋅ 2) = 24a ⋅ 12 6(4a ⋅ 2) = (4a ⋅ 2) ⋅ 6 6(4a ⋅ 2) = 6 ⋅ 4a ⋅ 2 6(4a ⋅ 2) = (6 ⋅ 4a) ⋅ 2 Which algebraic property could be used to rewrite 4x + 2y as 2y + 4x? Associative Property of Addition Associative Property of Multiplication

OpenStudy (petlovers123):

Commutative Property of Addition Commutative Property of Multiplication Which of the following demonstrates the Distributive Property? 3(4a + 2) = 12a + 6 3(4a + 2) = 12a + 2 3(4a + 2) = 4a + 6 3(4a + 2) = 7a + 5 Which of the following shows 2 + (x + 3y) rewritten using the Associative Property of Addition? 2 + x + 3y (2 + x) + 3y 2x + (3 + y) x + (2 + 3y)

OpenStudy (petlovers123):

@jhonyy9

jhonyy9 (jhonyy9):

do you know what mean eaxctly commutativ ?

OpenStudy (petlovers123):

Do i know what commutative property meas

OpenStudy (petlovers123):

means

jhonyy9 (jhonyy9):

indifferent where is 2 in front of parantheses or next parantheses the result will be the same

OpenStudy (petlovers123):

b.

OpenStudy (petlovers123):

@jhonyy9

jhonyy9 (jhonyy9):

not b. i think is a.

jhonyy9 (jhonyy9):

indifferent where is 2 in front of parantheses or next to parantheses

OpenStudy (petlovers123):

for the second question i think d.

jhonyy9 (jhonyy9):

yes d. is right for the second one sure

OpenStudy (petlovers123):

i dont know the third one